# Adaptive Layer-skipping in Pre-trained LLMs

### v2

@@ Abstract
Various layer-skipping methods have been proposed to accelerate token generation in large language models (LLMs). However, they have overlooked a fundamental question: How do computational demands vary across the generation of different tokens? In this work, we introduce FlexiDepth, a method that dynamically adjusts the number of Transformer layers used in text generation. By incorporating a plug-in router and adapter, FlexiDepth enables adaptive layer-skipping in LLMs without modifying their original parameters. Introducing FlexiDepth to Llama-3-8B model achieves layer skipping of 8 layers out of 32, and meanwhile maintains the full 100% benchmark performance. Experimental results with FlexiDepth demonstrate that computational demands in LLMs significantly vary based on token type. Specifically, generating repetitive tokens or fixed phrases requires fewer layers, whereas producing tokens involving computation or high uncertainty requires more layers. Interestingly, this adaptive allocation pattern aligns with human intuition. To advance research in this area, we open sourced FlexiDepth and a dataset documenting FlexiDepth’s layer allocation patterns for future exploration.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in various tasks, including translation [45], code generation [17], and math reasoning [6]. Currently, LLMs typically generate each token by performing a full forward pass through all Transformer decoder layers. However, such uniform allocation is counter-intuitive, as simpler tasks intuitively require fewer computational resources, whereas more complex tasks demand more extensive processing. This uniform allocation not only results in computational inefficiencies but may also contribute to overfitting in LLMs. To mitigate the inefficiencies of uniform allocation, various layer-skipping techniques have been proposed for acceleration. One line of research leverages statistical information, such as the difference between layer inputs and outputs, to identify and skip less important layers [27, 14, 44]. Another collection of methods involves early-exit [38, 19, 35], where a confidence measure or a router dynamically determines whether to bypass all subsequent layers at an intermediate point of LLM. While these techniques have successfully reduced computational costs, they overlook a fundamental question: How do computational demands vary across the generation of different tokens?
@@
Contrast to layer skipping methods requiring training LLM from scratch, i.e. mixture-of-depth (MoD) [32], FlexiDepth enables adaptive layer-skipping in pre-trained LLMs without modifying their original parameters. Specifically, FlexiDepth introduces two light-weight plug-in modules at each decoder layer: (1) a router that makes binary decisions about whether to skip the layer, and (2) a adapter that resolves representation mis-alignments caused by layer skipping. The router and adapter are the only trainable components within our framework. All other parameters are inherited from a pre-trained LLM and remain frozen. At each layer, the input hidden states are first evaluated by the router and directed to go through or skip the layer. For hidden states that go through the layer, they are processed by the original attention and FFN modules of the LLM. Conversely, hidden states that skip the layer bypass these modules and are instead processed by the adapter. The adapter’s role is to transform these skipped hidden states, aligning their representation to match those undergo full processing. Finally, both the adapted hidden states and the processed hidden states are combined to produce the layer output.
@@
Figure 1: DepthMap illustrating layer-skipping patterns when applying FlexiDepth to Llama-3-8B-Instruct. The light-to-dark blue gradient indicates layer usage ranging from 16 to 32.
To ensure compatibility with autoregressive generation, FlexiDepth computes the KV cache [29] for all hidden states, including those that skip this layer. This crucial step preserves the full context required for the attention mechanism. Without maintaining the KV cache, subsequent queries would be unable to attend to the hidden states that skipped the layer, rendering those tokens permanently unavailable for future generation. This would compromise the model’s ability to produce coherent and contextually accurate sequences. Preparing the full KV cache offers a simple method to maintain the integrity of autoregressive generation.
@@
We comprehensively evaluate FlexiDepth across a diverse set of benchmarks. When applied to a Llama-3-8B-Instruct model [9], FlexiDepth preserves full performance (100.7%) while skipping 8 out of 32 layers, significantly outperforming existing layer-skipping methods, particularly in continuous generation tasks.
@@
Beyond performance gain, FlexiDepth can reveal layer allocation patterns of pre-trained LLMs, addressing the question of how computational demands vary when generating different types of tokens. To demonstrate this phenomenon, we created a colored map that shows the number of layers used to generate each token, and call it "DepthMap". Figure 1 (left) shows that summarization typically requires more layers on average than extractive question answering. Similarly, Figure 1 (right) reveals that in mathematical reasoning tasks like addition, tokens on the left-hand side of equations need fewer layers to generate than those on the right-hand side. This type of adaptive allocation appears to resonate with human intuition. For instance, in language tasks, humans typically exert more effort in summarization to condense the entire context into a concise form. In contrast, extractive question answering often relies on retrieval and extraction, making it easier to identify the answer. Likewise, in mathematical tasks, deriving the right-hand side of an equation involves computation that consumes more steps. Shifting from token-specific patterns to broader layer usage, we observe that layer utilization follows a long-tail distribution, with early and final layers used more frequently than middle layers. To inspire further research along this direction, we compile a dataset documenting the layer allocation behaviors of FlexiDepth across various tasks.
@@
Figure 2: The FlexiDepth layer. Left: Full-processing path where hidden states undergo the pre-trained attention and FFN modules. Right: Skipping path where hidden states bypass the attention module and processed by a lightweight adapter. The router and adaptor (in red) are the only trainable components within the FlexiDepth Block.
@@ 2 Method
In this section, we present the detailed implementation of FlexiDepth, which consists of multiple FlexiDepth layers. Each FlexiDepth layer enhances a standard pre-trained decoder layer with a plug-in router and adapter, allowing adaptive layer-skipping without modifying the original parameters. Different from Coda [22] and COLT5 [1] that adopt adapters for conditional computation in transformer encoders, our method focuses on seamless integration with decoder-only models. As illustrated in Figure 2, in each FlexiDepth layer, the router first processes normalized input hidden states to make skipping decisions. Based on the router’s output, each hidden state either follows the full-processing path (Figure 2, left) or the skipping path (Figure 2, right). The outputs from both paths are then combined to produce the layer output. In the following parts, we elaborate on the core designs of FlexiDepth, including router design, attention skipping, FFN skipping, and layer-skipping loss.
2.1 Router design
FlexiDepth employs a router to compute gating scores for routing. Let $X=[x_{1},x_{2},\ldots,x_{T}]\in\mathbb{R}^{T\times d}$ denote the input hidden states, where $T$ is the sequence length and $d$ is the hidden dimension. The router computes their corresponding gating scores $G=[g_{1},g_{2},\ldots,g_{T}]\in\mathbb{R}^{T}$ as follows:
@@ 2.1 Router design
where Norm refers to the RMSNorm [43], and $\sigma$ is the sigmoid function to ensure $g_{i}\in(0,1)$. In previous works [32, 37], the router is typically implemented as a simple linear transformation, and jointly optimized with the transformer during training. However, in our setting, the original parameters of the pre-trained transformer are frozen. Merely relying on a linear transformation is insufficient to capture the nuances of the hidden states for routing. To address this issue, we design a parameter-efficient router based on a bottlenecked MLP:
@@
Upon obtaining the gating scores $G$, we implement a threshold-based routing mechanism with a predefined threshold $\tau$. For each hidden state $x_{i}$, if $g_{i}>\tau$, it will be routed to the full-processing path (Figure 2, left), with the output multiplied by $g_{i}$ to maintain gradient flow to the router. Conversely, if $g_{i}\leq\tau$, it will be routed to the skipping path (Figure 2, right), with the output multiplied by $(1-g_{i})$.
@@ 2.2 Attention skipping
In the skipping path of FlexiDepth (Figure 2, right), the input hidden states simply bypass the attention module via a shortcut. However, this straightforward approach can lead to a significant loss of contextual information, resulting in performance degradation. The attention mechanism relies on query, key, and value vectors to enable tokens to attend to each other, capturing long-range dependencies for generation. When hidden states skip the attention module, their corresponding key and value vectors are not generated, preventing subsequent tokens from attending to them. This issue is illustrated in Figure 3 (middle), where in this example, hidden states $x_{1}$ and $x_{3}$ are routed to skip the layer. Naively skipping the attention module for these tokens results in missing key-value pairs (i.e., $k_{1},v_{1}$ and $k_{3},v_{3}$). Therefore, the subsequent tokens, such as $q_{4}$, can no longer attend to them.
To address this issue, we propose a straightforward solution: we continue computing the keys and values for hidden states that skip the layer. Specifically, while we omit the query vectors (e.g., $q_{1},q_{3}$) and their associated scaled dot-product operations, we retain the calculations of the corresponding keys ($k_{1},k_{3}$) and values ($v_{1},v_{3}$). This approach, as demonstrated in Figure 3 (right), ensures that future tokens maintain access to the full set of contextual information, preserving the integrity of autoregressive generation.
@@ 2.3 FFN skipping
We observe that directly skipping the feedforward network (FFN) via a simple shortcut leads to significantly degrade model performance. Unlike the attention module, which only consists of linear $q$, $k$, $v$ transformations, the FFN module introduces nonlinearities. Consequently, hidden states transformed by the FFN might not share the same latent space as those skipping it. To mitigate this issue, we employ a lightweight adapter to align their representations. This adapter follows the same structure as the FFN but features a significantly reduced intermediate dimension, with a reduction factor of 16x. As shown in Figure 2 right, the adapter is placed at the same position as the skipped FFN.
@@ 3.1 Implementation details
We implement FlexiDepth on the pre-trained Llama-3-8B-Instruct model [9], which consists of 32 transformer layers. To enable adaptive layer-skipping, we convert the latter 16 layers into FlexiDepth layers, as prior works [27, 36] and empirical studies suggest that skipping earlier layers degrades performance. For each FlexiDepth layer, the router, as defined in Equation 3, uses a bottleneck dimension $d_{r}=\frac{1}{16}d$, where $d$ is the hidden dimension. The router’s gating function is implemented using SparseMixer [24, 25] to ensure differentiability. The adapter has the same structure as the original FFN but reduces the intermediate dimension by a factor of 16. For the loss function in Equation 4, we set the coefficient $\alpha=1\times 10^{-3}$, which results in approximately 8 layers skipped during generation. We train FlexiDepth on the Tulu-v2 dataset [18] for 3 epochs using the AdamW [26] optimizer with a learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$. We use a warmup ratio of 0.03 and a global batch size of 64. The training takes approximately 7 hours on 8 NVIDIA A100-PCIE-40GB GPUs.
@@ 3.2 Main results
In this part, we present the experimental results of applying the proposed FlexiDepth to the pre-trained Llama-3-8B-Instruct model [9].
@@
We evaluate FlexiDepth across a diverse set of tasks. For single-token generation benchmarks, we include: MMLU [15], HellaSwag [41], and Winogrande [34]. For multi-token generation tasks, we test on: GSM8K [5], HumanEval [4], CoQA [33]. Evaluations are conducted using the lm-evaluation-harness toolkit [11], with 5-shot settings for MMLU, HellaSwag, Winogrande, and GSM8K, and zero-shot settings for HumanEval and CoQA. The evaluation metrics are accuracy (acc) for MMLU, normalized accuracy (acc_norm) for HellaSwag, accuracy (acc) for Winogrande, exact-match (EM) for GSM8K, pass-at-1 (Pass@1) for HumanEval, and F1 score (F1) for CoQA.
@@
We compared FlexiDepth with previous layer-skipping methods that are compatible with LLMs. All baseline methods are applied to the LLama-3-8B-Instruct model, which consists of 32 transformer layers. We define $k$ as the number of layers to skip during generation, with all methods configured to skip the same number of layers for fair comparison. The baseline methods are as follows: (1) LayerSkip [10]: This early-exit method skips the last $k$ consecutive layers during decoding and employs speculative decoding [23] to refine generation results. For comparison, we disable speculative decoding. (2) ShortGPT [27]: This approach prunes $k$ layers deemed less important by assessing their input-output difference. (3) LaCo [40]: This method employs a layer-collapse strategy to reduce the model by $k$ layers, merging subsequent layers into a prior one using a Reserving-Differences-while-Seeking-Common (RDSC) approach. Please read the original paper for the details. (4) MindSkip [13]: This method explores attention, FFN, and layer skipping via a simple linear router, finding that skipping FFN or entire layers leads to significant performance degradation. For comparison, we employ its layer skipping setting and train it on the same tulu-v2 [18] dataset for comparison.

### v3

@@ Abstract
Various layer-skipping methods have been proposed to accelerate token generation in large language models (LLMs). However, limited attention has been paid to a fundamental question: How do computational demands vary across the generation of different tokens? In this work, we introduce FlexiDepth, a method that dynamically adjusts the number of Transformer layers used in text generation. By incorporating a plug-in router and adapter, FlexiDepth enables adaptive computation in LLMs without modifying their original parameters. Applied to Llama-3-8B, it skips 8 out of 32 layers while maintaining full benchmark performance. Our experiments reveal that computational demands in LLMs significantly vary based on token type. Specifically, generating repetitive tokens or fixed phrases requires fewer layers, whereas producing tokens involving computation or high uncertainty requires more layers. Despite the computational savings, FlexiDepth does not yet achieve wall-clock speedup due to varied skipping patterns and I/O overhead. To inspire future work and advance research on practical speedup, we open-sourced FlexiDepth and a dataset documenting its layer allocation patterns.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in various tasks, including translation (Zhu et al., 2024), code generation (Hui et al., 2024), and math reasoning (DeepSeek-AI et al., 2025). Currently, LLMs typically generate each token by performing a full forward pass through all Transformer decoder layers. However, such a uniform allocation is counter-intuitive, as simpler tasks intuitively require fewer computational resources, whereas complex tasks demand more processing. This uniform allocation not only results in computational inefficiencies but may also contribute to overfitting in LLMs. In response to these challenges, various layer-skipping techniques have been proposed. One line of research leverages statistical information, such as the difference between layer inputs and outputs, to identify and skip less important layers (Men et al., 2024; He et al., 2024b; Zhang et al., 2024). Another collection of methods involves early-exit (Varshney et al., 2024; Jazbec et al., 2024; Schuster et al., 2022), where a confidence measure or a router dynamically determines whether to bypass all subsequent layers at an intermediate point of LLM. While these techniques have successfully reduced computational costs, limited attention has been paid to a fundamental question: How do computational demands vary across the generation of different tokens?
@@
Contrast to layer skipping methods requiring training LLM from scratch, i.e. mixture-of-depth (MoD) (Raposo et al., 2024), FlexiDepth enables adaptive layer-skipping in pre-trained LLMs without modifying their original parameters. Specifically, FlexiDepth introduces two light-weight plug-in modules at each decoder layer: (1) a router that makes binary decisions about whether to skip the layer, and (2) a adapter that resolves representation mis-alignments caused by layer skipping. The router and adapter are the only trainable components within our framework. All other parameters are inherited from a pre-trained LLM and remain frozen. At each layer, the input hidden states are first evaluated by the router and directed to go through or skip the layer. For hidden states that go through the layer, they are processed by the original attention and FFN modules of the LLM. Conversely, hidden states that skip the layer are instead processed by the adapter. The adapter’s role is to transform these skipped hidden states, aligning their representation to match those that undergo full processing. Finally, both the processed hidden states and the adapted hidden states are combined to produce the layer output. To optimize the router and adapter, we introduce a layer skipping loss that works in conjunction with the original next-token prediction loss. This loss function encourages layer skipping by penalizing the number of layers used for generation, while maintaining generation quality through the prediction loss. After training, the model can dynamically adjust its layer usage based on the input.
@@
Figure 1: DepthMap illustrating layer-skipping patterns when applying FlexiDepth to Llama-3-8B-Instruct. The light-to-dark blue gradient indicates layer usage ranging from 16 layers to 32 layers.
@@
We comprehensively evaluate FlexiDepth across a diverse set of benchmarks. When applied to Llama-3-8B-Instruct (Dubey et al., 2024), FlexiDepth preserves full performance (100.7%) while skipping 8 out of 32 layers, significantly outperforming existing layer-skipping methods, particularly in continuous generation tasks.
@@
FlexiDepth also reveals how computational demands vary when generating different types of tokens. For demonstration, we created a colored map that shows the number of layers used to generate each token, which we call ’DepthMap’. Figure 1 (left) shows that summarization typically requires more layers on average than extractive question answering. Similarly, Figure 1 (right) reveals that in mathematical reasoning tasks like addition, tokens on the left-hand side of equations need fewer layers to generate than those on the right-hand side. This type of adaptive allocation appears to resonate with human intuition. For instance, humans typically exert more effort in summarization tasks to condense the entire context into a concise form, while finding extractive question answering easier as it only requires retrieval and extraction. Likewise, in mathematical tasks, deriving the right-hand side of an equation involves computation that consumes more steps. Shifting from token-specific patterns to overall layer usage, we observe that layer utilization follows a long-tail distribution, with early and final layers used more frequently than middle layers. To inspire further research along this direction, we compile a dataset documenting the layer allocation behaviors of FlexiDepth across various tasks.
@@
Figure 2: The FlexiDepth layer. Left: Full-processing path where hidden states undergo the pre-trained attention and FFN modules. Right: Skipping path where hidden states bypass the attention module and processed by a lightweight adapter. The router and adaptor (in red) are the only trainable components.
@@ 2 Method
In this section, we introduce FlexiDepth, a framework for adaptive layer-skipping in pre-trained LLMs. Each FlexiDepth layer augments a standard decoder layer with a lightweight router and adapter, keeping the original parameters frozen. The router evaluates normalized input hidden states to compute the gating score $g$. Based on a threshold $\tau$, hidden states with $g>\tau$ are routed to the full-processing path (Figure 2, left), where they undergo the pre-trained attention and FFN modules. Hidden states with $g\leq\tau$ are routed to the skipping path (Figure 2, right), where they are instead processed by a lightweight adapter. In the following parts, we elaborate on the core designs of FlexiDepth, including routing, attention skipping, FFN skipping, and layer-skipping loss.
2.1 Routing
FlexiDepth employs a router at each layer to compute gating scores for routing. Let $X=[x_{1},x_{2},\ldots,x_{T}]\in\mathbb{R}^{T\times d}$ denote the input hidden states of the current layer, where $T$ is the sequence length and $d$ is the hidden dimension. The router computes their corresponding gating scores $G=[g_{1},g_{2},\ldots,g_{T}]\in\mathbb{R}^{T}$ as follows:
@@ 2.1 Routing
where Norm refers to the RMSNorm (Zhang & Sennrich, 2019), and $\sigma$ is the sigmoid function to ensure $g_{i}\in(0,1)$. In previous works (Raposo et al., 2024; Tan et al., 2024), the router is typically implemented as a simple linear transformation, and jointly optimized with the transformer during training. However, in our setting, the original parameters of the pre-trained transformer are frozen. Relying on a linear transformation is insufficient to capture the nuances of the hidden states for routing. To address this issue, we design a parameter-efficient router based on a bottlenecked MLP:
@@
After computing the gating scores $G$, we apply a threshold-based routing mechanism using a predefined threshold $\tau$. For each hidden state $x_{i}$, if $g_{i}>\tau$, it is routed to the full-processing path (Figure 2, left), where the output is scaled by $g_{i}$ to ensure gradient flow to the router. If $g_{i}\leq\tau$, the hidden state is routed to the skipping path (Figure 2, right), with the output scaled by $(1-g_{i})$. This process is formalized as:
$x_{i}^{\prime}=\begin{cases}g_{i}\cdot\text{FFN}(\text{Norm}(\text{Attn}(\text{Norm}(x_{i}))+x_{i}))+\text{Attn}(\text{Norm}(x_{i}))+x_{i},&\text{if }g_{i}>\tau\\ (1-g_{i})\cdot\text{Adapter}(\text{Norm}(x_{i}))+x_{i},&\text{if }g_{i}\leq\tau\end{cases}$ (3)
where $x_{i}^{\prime}$ is the output hidden state for token $i$, Norm represents layer normalization, FFN denotes the feed-forward network, and $\text{Adapter}(\cdot)$ is the lightweight adapter.
@@ 2.2 Attention skipping
In the skipping path of FlexiDepth (Figure 2, right), the input hidden states simply bypass the attention module via a shortcut. However, this straightforward approach can lead to a significant loss of contextual information, resulting in performance degradation. In autoregressive generation, the attention mechanism computes query, key, and value vectors for each token, where keys and values are cached (KV Cache) (Pope et al., 2023) to avoid redundant computation in subsequent decoding steps. When hidden states skip the attention module, their corresponding key and value vectors are not generated, preventing subsequent tokens from attending to them. This issue is illustrated in Figure 3 (middle), where in this example, hidden states $x_{1}$ and $x_{3}$ are routed to skip the layer. Naively skipping the attention module for these tokens results in missing key-value pairs (i.e., $k_{1},v_{1}$ and $k_{3},v_{3}$). Therefore, the subsequent tokens, such as $q_{4}$, can no longer attend to them.
To address this issue, we propose a straightforward solution: we continue computing the keys and values for hidden states that skip the layer. Specifically, while we can omit the query vectors (e.g., $q_{1},q_{3}$) and their associated scaled dot-product operations, we retain the calculations of the corresponding keys ($k_{1},k_{3}$) and values ($v_{1},v_{3}$). This approach, as demonstrated in Figure 3 (right), ensures that future tokens maintain access to the full set of contextual information, preserving the integrity of autoregressive generation.
@@ 2.3 FFN skipping
We observe that directly skipping the feedforward network (FFN) via a simple shortcut significantly degrades model performance. Unlike the attention module, which only involves linear transformations, the FFN module introduces nonlinearities. Consequently, hidden states transformed by the FFN may not share the same latent space as those that skip it. To mitigate this issue, we employ a lightweight adapter to align their representations. This adapter follows the same structure as the FFN but features a significantly reduced intermediate dimension. As shown in Figure 2 (right), the adapter is placed at the same position as the skipped FFN.
@@ 3.1 Implementation details
We implement FlexiDepth on the pre-trained Llama-3-8B-Instruct model (Dubey et al., 2024), which consists of 32 transformer layers. To enable adaptive layer-skipping, we convert the latter 16 layers into FlexiDepth layers, as prior works (Men et al., 2024; Sun et al., 2024) and empirical studies suggest that skipping earlier layers degrades performance. For each FlexiDepth layer, the router, as defined in Equation 2, uses a bottleneck dimension $d_{r}=\frac{1}{16}d$, where $d$ is the hidden dimension. The router’s gating function is implemented using SparseMixer (Liu et al., 2023a; b) to ensure differentiability. The adapter has the same structure as the original FFN but reduces the intermediate dimension by a factor of 16. For the loss function in Equation 5, we set the coefficient $\alpha=1\times 10^{-3}$, which results in approximately 8 layers skipped during generation. We train FlexiDepth on the Tulu-v2 dataset (Ivison et al., 2023) for 3 epochs using the AdamW (Loshchilov & Hutter, 2019) optimizer with a learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$. We use a warmup ratio of 0.03 and a global batch size of 64. The training takes approximately 7 hours on 8 NVIDIA A100-PCIE-40GB GPUs.
@@ 3.2 Main results
In this part, we present the experimental results of applying the proposed FlexiDepth to the pre-trained Llama-3-8B-Instruct model (Dubey et al., 2024).
@@
We evaluate FlexiDepth across a diverse set of tasks. For single-token generation benchmarks, we include: MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), and Winogrande (Sakaguchi et al., 2020). For multi-token generation tasks, we test on: GSM8K (Cobbe et al., 2021), HumanEval (Chen et al., 2021), CoQA (Reddy et al., 2019). Evaluations are conducted using the lm-evaluation-harness toolkit (Gao et al., 2024), with 5-shot settings for MMLU, HellaSwag, Winogrande, and GSM8K, and zero-shot settings for HumanEval and CoQA. The evaluation metrics are accuracy (acc) for MMLU, normalized accuracy (acc_norm) for HellaSwag, accuracy (acc) for Winogrande, exact-match (EM) for GSM8K, pass-at-1 (Pass@1) for HumanEval, and F1 score (F1) for CoQA.
@@
We compared FlexiDepth with previous layer-skipping methods that are compatible with LLMs. All baseline methods are applied to the LLama-3-8B-Instruct model, which consists of 32 transformer layers. We define $k$ as the number of layers to skip during generation, with all methods configured to skip the same number of layers for fair comparison. The baseline methods are as follows: (1) LayerSkip (Elhoushi et al., 2024): This early-exit method skips the last $k$ consecutive layers during decoding and employs speculative decoding (Leviathan et al., 2023) to refine generation results. For comparison, we disable speculative decoding. (2) ShortGPT (Men et al., 2024): This approach prunes $k$ layers deemed less important by assessing their input-output difference. (3) LaCo (Yang et al., 2024b): This method employs a layer-collapse strategy to reduce the model by $k$ layers, merging subsequent layers into a prior one using a Reserving-Differences-while-Seeking-Common (RDSC) approach. (4) MindSkip (He et al., 2024a): This method explores attention, FFN, and layer skipping via a simple linear router, finding that skipping FFN or entire layers leads to significant performance degradation. For comparison, we employ its layer skipping setting and train it on the same tulu-v2 (Ivison et al., 2023) dataset.
@@
Model MMLU Hellaswag Winogrande GSM8K HumanEval CoQA Retain %
Llama-2-13B 0.492 0.727 0.714 0.236 0.102 0.790 100.0%
FlexiDepth 0.481 0.741 0.716 0.253 0.096 0.780 100.2%
Skipped 7.08 1.98 5.00 6.65 11.61 6.19 -
Llama-3-8B 0.673 0.706 0.744 0.679 0.299 0.784 100.0%
FlexiDepth 0.663 0.743 0.756 0.657 0.323 0.803 102.1%
Skipped 4.12 2.00 3.97 10.42 9.55 7.44 -
Qwen-2.5-3B 0.651 0.702 0.639 0.576 0.487 0.705 100.0%
FlexiDepth 0.643 0.701 0.679 0.583 0.476 0.741 101.5%
Skipped 1.23 1.25 1.15 1.71 1.69 1.88 -